Wilma is 13 years older than Bev. In nine years, Wilma will be twice as old as Bev. How old is Bev?
Found 2 solutions by ikleyn, rothauserc:
Answer by ikleyn(52781)

.
In nine years, Bew will be 13, while Wilma will be 26.


Hence, now Bew is 13-9 = 4 years old,  while Wilma is 26-9 = 17.

------------
There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.


Answer by rothauserc(4718)

let w be Wilma's age and b be Bev's age
:
1) w = b +13
:
2) w +9 = 2(b+9)
:
use equation 1 and substitute for w in equation 2
:
b +13 +9 = 2b +18
:
b = 4
w = 4 +13 = 17
:
*****************************
Wilma is 17 and Bev is 4
:
check answer in equation 2
:
17 +9 = 2(4+9)
:
26 = 26
********************************
: